40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log, Page 234 of 327  Not logged in ELOG logo
ID Date Author Type Category Subject
  4736   Wed May 18 07:13:00 2011 SureshUpdateRF SystemDemod board measurements

I measured the amplitude and phase imbalances of the demod boards which have been modified.  This is just a basic health check.  We hope to use the script that Kiwamu is developing for a more accurate test.  The script can also use these measurements as a sanity check.  POP110  requires some further attention. 

Demod_Board_measurements.png

 

The RF distribution box outputs corresponding to the demod board (eg. AS55_LO --> AS55_demod) were used as LO sources.  The RF signal was generated with a Marconi and held a kHz away from the LO frequency.  The amplitude and phase unbalance were measured with SR785.  The RF Power meter was used to check the LO power in each case.

 

 

Attachment 1: Demod_Board_measurements.png
Demod_Board_measurements.png
  4735   Tue May 17 22:50:00 2011 JamieConfigurationLSCNew digital lockin added to LSC model/screen

I added a lockin to the LSC model, and added it's corresponding control to the LSC screen.  Here's is what I added to the LSC model:

lsc-mdl-lockin.png

On the left are the froms from the normalized RF PD outputs. Those go into a matrix, whose single row goes into the lockin signal input. The clock output from the lockin goes into the LSC output matrix (last row).

Here is what I added to the LSC master screen:

lsc-screen-lockin.png

I also modified the lockin overview screen:

lsc-lockin-screen.png

I think this new screen looks a lot cleaner.  Maybe we could start using this one as a template for lockin screens.

  4734   Tue May 17 19:38:32 2011 kiwamuUpdateSUSwrong connection on 1X5

Today Steve was working around the 1X5 rack to strain relief the cable jungles and the jungle is now getting less jungle.

During the work he disconnected and reconnected some cables.

So for a doublecheck I checked all the suspensions to see if the suspensions are still healthy or not.

Aha, then I found a mistake.

 

See the pictures below. It's a very subtle difference. This wrong connection prevented MC1 and MC3 from damping.

wrong.png correct.png

  4733   Tue May 17 18:09:13 2011 Jamie, KiwamuConfigurationCDSc1sus and c1auxey crashed, rebooted

c1sus and c1auxey crashed, required hard reboot

For some reason, we found that c1sus and c1auxey were completely unresponsive.  We went out and gave them a hard reset, which brought them back up with no problems.

This appears to be related to a very similar problem report by Kiwamu just a couple of days ago, where c1lsc crashed after editing the C1LSC.ini and restarting the daqd process, which is exactly what I just did (see my previous log).  What could be causing this?

  4732   Tue May 17 17:01:22 2011 jamieConfigurationCDSUpdate LSC channels from _DAQ to _DQ

As of RCG version 2.1, recorded channels use the suffix "_DQ", instead of "_DAQ".  I just rebuilt and installed the c1lsc model, which changed the channel names, therefore hosing the old daq channel ini file.  Here's what I did, and how I fixed it:

$ ssh c1lsc
$ cd /opt/rtcds/caltech/c1/core/trunk
$ make c1lsc
$ make install-c1lsc
$ cd /opt/rtcds/caltech/c1/scripts
$ ./startc1lsc
$ cd /opt/rtcds/caltech/c1/chans/daq
$ cat archive/C1LSC_110517_152411.ini | sed "s/_DAQ/_DQ/g" >C1LSC.ini
$ telnet fb 8087
daqd> shutdown

 

  4731   Tue May 17 16:53:22 2011 kiwamuUpdateSUSoplev sign was wrong on ETMY

Last night I found that the sign of the oplev control of PITCH on ETMY was wrong. I flipped it to the correct sign.

We've been locking the Y arm by feeding a signal back to ITMY  because pushing ETMY somehow made the lock unstable in the angular motion.

After the correction of the oplev contol sign, I was able to keep the lock robustly by pushing ETMY.

  4730   Tue May 17 11:45:20 2011 JamieConfigurationCDSpurged non-c1 site files from rtcds checkout of cds_user_apps
I purged all of the working copy checkouts of site files for all sites that are not c1 from the rtcds cds_user_apps working directory (/opt/rtcds/caltech/c1/userapps/trunk). I first checked that there were no outstanding changes, and then did the following (in bash):
cd /opt/rtcds/caltech/c1/userapps
rm -rf trunk
svn update --depth=files trunk
svn update --depth=empty trunk/{CDS,ISI,ISC,PSL,SUS}
svn update trunk/{CDS,ISI,ISC,PSL,SUS}/{c1,common}
  4729   Tue May 17 01:05:56 2011 kiwamuUpdateLSCAsymmetry measurement prep : recentering works

I re-centered beams on several PDs and a camera including :

  AS55, ETMY_QPD, TRY and ETMYT_CCD.

 

The most important one was AS55.

When I was locking each arm I found that the error signal from AS55 was very coupled to the angular motion of the arms.

I checked the beam on the AS55 RFPD and found the beam on the edge of the photo diode. This is possibly because Valera and I had been touching the input beam alignment.

At that time the DC signal from AS55 without aligning PRM and SRM was about 5 mV.

Adjusting the beam position by a steering mirror brought the DC signal up to 20 mV.

Then the lock of each arm became more stable.

  4728   Mon May 16 17:11:28 2011 steveUpdateElectronicsstrain relieved 1X1

1X1 is strain relieved in the back. I will use similar approach on the rest of the racks .

Attachment 1: P1070724.JPG
P1070724.JPG
  4727   Mon May 16 15:46:05 2011 KojiUpdateGeneraloptical layout viewable on the online version of autocad

I have made my own AutoCAD WS account and put the latest 40m layout.
AutoCAD WS is a free cloud service which enable us to browse/edit the DWG files.

You can view/play/print it from the following link even without making the account.

https://www.autocadws.com/main/publish?link=RVFmNEtOd3EzNVFtRXcx

If you make your account you can actually save it although the editing capability is somewhat limited.

Notes:
- It needs the Flash plug-in and Internet connection.
- I think Safari is the most stable platform on Mac. (i.e. I had some malfunctions with Firefox and Chrome)

  4726   Mon May 16 11:47:59 2011 kiwamuUpdateASSc1ass update part II

The medm screen for c1ass started being modified to be more user-friendly.

The modification is still ongoing, but the goal is to make a screen which anyone can easily understand and play with.

 

Still to do : ( need a volunteer )

- Modification of the screens

- Commission the input beam and X-arm servos

- Make scripts for X-arm

- Measure the PZT mirrors' matrix for the translation and angle

c1ass.png

Quote from #4709

Here the status of the dither alignment or c1ass:

Still to do:

- Commission the input beam and X-arm servos

- Make scripts for X-arm

 

  4725   Mon May 16 11:14:36 2011 ranaUpdateIOOwiring diagram for IP-POS

Just FYI, the 3113 is a 12-bit ADC, so we won't get very good resolution out of these. We should get one of those purple boxes.

Also, I corresponded some with Dave Barker. The 40m Wiki at LHO is down because of disk errors. He's working on it and will let us know. But suggests that we move it to Caltech since he'll turn the box off at the end of the year.

Mon May 16 13:57:49 2011    Wiki is back up.

  4724   Mon May 16 10:05:02 2011 kiwamuUpdatePhotosRe:ETMY optical bench

You are right. We should change or rotate the mirror mount.

Actually when Suresh and I were putting the mirror we rotated the mount  by 90 deg such that the fat side of the mount is at left had side.

It was because the fat side had been clipping the oplev beam when the fat side is at right.

At that moment we were blocking the green beam to only see the faint IR beam with a sensor card, so we haven't checked the green beam.

Anyway the mount is apparently not good for the green beam.

Quote from #4723

I didn't notice it the other day when I was working on putting in the trans QPD, but do we need to switch the mirror mount for the first turning mirror of the IR trans beam, which the green transmits through to go into the cavity?  It seems like we've set ourselves up for potential clipping.

 

  4723   Sun May 15 21:27:51 2011 JenneUpdatePhotosETMY optical bench

I didn't notice it the other day when I was working on putting in the trans QPD, but do we need to switch the mirror mount for the first turning mirror of the IR trans beam, which the green transmits through to go into the cavity?  It seems like we've set ourselves up for potential clipping.

Quote:

Just for a record. This is the latest picture of the ETMY optical bench.

I will upload this picture on the wiki after the wiki gets up.

ETMY_ss.jpg

 

  4722   Sun May 15 19:55:15 2011 kiwamuUpdatePhotosETMY optical bench

Just for a record. This is the latest picture of the ETMY optical bench.

I will upload this picture on the wiki after the wiki gets up.

ETMY_ss.jpg

  4721   Sun May 15 19:10:12 2011 kiwamuUpdateLSCCalibration of actuators : BS, ITMX and ITMY

The AC response of the actuators on BS, ITMX and ITMY were re-measured by another technique.

Last time I estimated them by measuring the open-loop transfer functions, but this time the responses were measured in a more direct way.

The measured AC responses (60 Hz - 200 Hz) are :

      BS   = 1.643e-98 / f2  [m/counts] (corrected based on the plot below - Manasa)

     ITMX = 3.568e-9 / f2 [m/counts]

     ITMY = 3.542e-9 / f2 [m/counts]

Next : measurement of the PRM actuator response


(The technique) 

 This time a technique that Rana told me a week ago was used.

This technique allows us to directly measure the response of an actuator at high frequency without any loop corrections.

First of all, MICH has to be locked to keep MICH within the linear range of the error signal. So now MICH is a linear sensor to the mirror motions.

In the MICH control a steep low pass filter should be inserted in order to avoid unwanted effects from the control loop at the high frequencies.

For example I put a low pass filter composed of an elliptical filter whose cut-off frequency is at 50 Hz such that the control loop doesn't push the mirrors above the cut-off frequency.

Hence the error signal of MICH above 50 Hz directly corresponds to the motion of the mirrors including BS, ITMX and ITMY.

Taking a transfer function from an actuator to the MICH error signal directly gives the actuator response.

In my measurements MICH was locked by feeding the signal back to BS. The plot below is the expected open-loop transfer function for the MICH control.

oltf.png

You can see that the open loop TF suddenly drops above 50 Hz. The UGF was at about 20 Hz, confirmed by looking at the loop oscillation on DTT.

 

(Measurement)

 In the technique the error signal has to be calibrated to [m]. This time AS55_Q was used and calibrated based on a peak-to-peak measurement.

The peak to peak value in the MICH error signal was 8 counts, which corresponds to the sensor efficiency of 4.72e+07 [counts/m].

Then I took transfer functions from each suspension (i.e. C1:SUS-XXX_LSC_EXC) to the error signal at AS55_Q over a frequency range from 60 Hz to 200Hz.

For the transfer function measurements I ran the swept sine on DTT to get the data. Note that the PD whitening filters were on.

The plot below is the results of the measurements together with the fitting lines.

calib_actuators.png

In the fitting I excluded the data pints at 60 Hz, because their coherence was low due to the power line noise.

  4720   Sun May 15 14:01:56 2011 kiwamuUpdateIOOwiring diagram for IP-POS

Here is a wiring diagram which shows how IP-POS (new official name is IB_POS) is connected.

Another thing we have to remember is : at some point we will also connect some more QPDs (e.g. POX, POY, AS, REFL and POP) in the same way.

They will also be acquired by the same slow machine : c1iscaux.

ip_pos.png

Quote from #4704

IP_POS is back. 

Hopefully I will make a simple wiring diagram such that we will never forget the connections.

  4719   Sun May 15 12:42:29 2011 kiwamuUpdateSUSf2p ratio adjutment done for all the suspensions

The f2p adjustment for all the suspensions are done (except for MC1,2,3)

Attachment 1: f2p_summary.pdf
f2p_summary.pdf f2p_summary.pdf f2p_summary.pdf f2p_summary.pdf f2p_summary.pdf f2p_summary.pdf f2p_summary.pdf
  4718   Sun May 15 03:58:19 2011 ranaUpdateCDSdiagonalization of MC input matrix

There has been some input matrix diagonalization in the past by Yuta and Kiwamu, but I find the automation to be not totally satisfactory.

It would be better if we could automatically fit the data to find the Suspended optic eigenfrequencies and then use that to get the matrix. So I wrote a peak fitter to get the matrix.

It gets the data from mafalda with NDS2, then it makes the PSDs, and then starts with some initial guesses (based on looking at the plots) and them uses fminsearch to get the peak frequencies and Q's.

Using the output of this, we can use Yuta's method and take the passive transfer functions with the free swing data (from April 30, so we got do do it quick) to get the input matrix.

 

Doing the SUS input matrix is nice for having good damping (as long as we remember to include SIDE), but my motivation is to produce a good null stream from the 4 face sensors so that we can estimate the sensor noises at all times.

Attachment 1: mc1.png
mc1.png
  4717   Sat May 14 14:50:21 2011 KojiUpdateLSC1Y2 5V Blown Fuse found -> Fixed

Incidentally, a blown fuse on 5V line at 1Y2 rack was found during the intallation of Sorensens.
The fuse (5A 125V) has been replaced and fixed.

When I plugged the fuse in, I heard some sound like relays were switched. Are there any relays in the LSC rack?

It was a 9th fuse from the top as seen in the picture.

Attachment 1: 1Y2_5V_Fuse.JPG
1Y2_5V_Fuse.JPG
  4716   Sat May 14 14:12:16 2011 KojiUpdateRF SystemDC power supply on RF distribution box has been replaced.

Key points of the power supply installation

  • We followed the grounding configuration for KEPCO except for the signal ground connection
  • AC power supply has been obtained from the local power strip. This also provides chassis earthing (for safety)
  • The chassis is connected to the shieldin of the DC supply cable. The other end should be isolated.
  • The low voltage side of Sorensen's DC outputs are connected in order to share the same reference  level.
  • The ground level is provided from the cross connect. The cable is connected between the cross connect ground to the sorencen.
    Unlike the KEPCO case, this cable does not have the current return, but just is to define the voltage level of those Sorensens.
  • New AC&DC cables have been nicely strain-relieved.

Quote:

[Steve, Koji, Suresh]

   We shifted two Sorensen power supplies from the Auxiliary rack next to 1X2 to 1Y2.  And have installed them there (pic below).  The local ground reference was picked up from the racks ground reference.  A shielded cable with two twisted pairs was used to make a new power cable for the RF rack.  Since we are using three of the four conductors (+18,+28 and ground), one of them is not connected to anything.  This situation can be improved in a future iteration when, for example, we might wish to relocate the Sorensens to a different rack.

   We are still working on changing the power supply to the RF source.  Will complete this early next week

 

Attachment 1: sorensen.png
sorensen.png
  4715   Fri May 13 23:04:58 2011 SureshUpdateRF SystemDC power supply on RF distribution box has been replaced.

[Steve, Koji, Suresh]

   We shifted two Sorensen power supplies from the Auxiliary rack next to 1X2 to 1Y2.  And have installed them there (pic below).  The local ground reference was picked up from the racks ground reference.  A shielded cable with two twisted pairs was used to make a new power cable for the RF rack.  Since we are using three of the four conductors (+18,+28 and ground), one of them is not connected to anything.  This situation can be improved in a future iteration when, for example, we might wish to relocate the Sorensens to a different rack.

   We are still working on changing the power supply to the RF source.  Will complete this early next week

P5130120.JPG

  4714   Fri May 13 22:45:37 2011 SureshUpdateRF SystemThe full set of 8 Demod boards is ready for testing

We have Completed the hardware changes to the full set of 8 demod boards.  The last one completed today is AS11.  I have collected the info on all the demod boards available so far in the table below.  As we measure the actual phase and amplitude unbalance we will expand this table to include new info.

The set of 8 demod boards
Demod Board S. No. Power Splitter Frequency range of splitter (MHz) Phase unbalance from datasheet (deg) Amplitude Unbalance from datasheet (dB)
AS11 121 PSCQ-2-51W 5 to 50 87.49 0.1
REFL11 021 PSCQ-2-51W 5 to 50 87.49 0.1
POY11 119 PSCQ-2-32 3.2 to 32 87.58 0.05
AS55 029 PSCQ-2-51W 5 to 50 no info no info
REFL55 118 PQW-2-90 30 to 90 90.14 0.92
POY11 119 PSCQ-2-32 3.2 to 32 87.58 0.05
POY22 020 PSCQ-2-32 3.2 to 32 90.26 0.02
POY110 120 PSCQ--120 80 to 120 90.88 0.58

 

  4713   Fri May 13 17:20:48 2011 Leo SingerConfigurationSUSTuned bounce and roll mode of ETMY suspension

I tuned the bounce and roll mode bandstops for ETMY, although it was difficult for me to tell if there was improvement with the bandstops on relative to the bandstops off because it seemed like the bounce and roll modes were being excited intermittently.  I'll take spectra with the filters both on and off during an evening next week.

  4712   Fri May 13 14:54:20 2011 Leo SingerUpdateComputersDiaggui fixed on pianosa

I fixed diaggui on pianosa.  Previously, it was not able to start because it depended on libreadline5, whereas Ubuntu distributed libreadline6.  Now pianosa has both libreadline5 and libreadline6, so diaggui works.

  4711   Fri May 13 01:51:56 2011 SureshUpdateRF SystemRF Status update

I have posted the attached RF status update and 1Y2 rack layout to the svn.

Attachment 1: 1Y2_Rack_Layout.pdf
1Y2_Rack_Layout.pdf
Attachment 2: RF_Work_Status.pdf
RF_Work_Status.pdf
  4710   Fri May 13 01:42:35 2011 kiwamuUpdateLSCready for Schnupp asymmetry measurement

[Valera / Kiwamu]

 We are able to lock each arm smoothly.  It is ready for the Schnupp asymmetry measurement.

( to be done )

 - Manual D-phase adjustment of the AS55 channel.

 - A script to adjust the D-phase.

 - Trigger logic for the boost filters.

 - Modification of some old alignment scripts to adopt them to the new LSC channels

  4709   Fri May 13 00:39:53 2011 valeraUpdateASSc1ass update

Here the status of the dither alignment or c1ass:

- Both pitch and yaw centering on ETMY/ITMY were closed simultatenously with ugf of ~1/30 Hz.

- I made a medm screen with beam positions as measured by the dither system.The snapshot is attached. There are visual perimeter alarms (red box around the display) to warn about arm power being low or the dither lines not being on. The screen has a pull down menu with 4 scripts:

. assUp - sets up the gains, phases and matricies for the dither system (both the spot centering and the input beam alignment)

. assOn - turns on the dithers and servo - just the Y-arm centering part at the moment

. assOff - turns off the servo and dither lines

. assDitherOn - turns on the dither lines but does not turn on the servo

- All scripts are in scripts/ASS and the medm screen is in medm/c1ass/master/

 

Still to do:

- Commission the input beam and X-arm servos

- Make scripts for X-arm

Attachment 1: c1assqpds.jpg
c1assqpds.jpg
  4708   Thu May 12 23:50:10 2011 SureshUpdateRF SystemPOY55_Demod board Hardware change completed

The Demod board with S. No. 022 (being used earlier as REFL11) has been modified.  It now has SCLF-65 as its input LP filter on the PD input line and a PQW-2-90 power splitter.  The unit functioning okay (I and Q signals are 90 deg apart.

The loss of Q output was traced to a possible loose solder joint and we now have both the I and Q signals after resoldering all components in the vicinity of U7 (Ref Schematic of D990511)

There is a strong oscillation around 350Hz present on I and Q signals of both REFL55_Demod and POY55_Demod.  Don't know the source. 

We have run out of power splitters to continue with the Demod board modification. We do not currently have an AS11_Demod board.  All the others are in place and ready for the I<->Q phase angle measurement.

In summary we now have the following Demod boards in place:

[ REFL11, POY11, REFL55, AS55, POY55, POY22, POY110]_Demod

 

  4707   Thu May 12 23:41:47 2011 ranaUpdateCDSMEDM Snapshots not working

Looks like 2 different MEDM Snapshot functiions (at least) are broken.

The regular update of the screens here as well as the usual "Update Snapshot" and view "previous snapshot" button on all of the auto-generated screens.

Also, how do we add the snapshot button to the custom made screens?

  4706   Thu May 12 23:12:40 2011 kiwamuUpdateCDSc1lsc crashed

This is my third time to crash a real-time machine. This time I crashed c1lsc.

I physically rebooted c1lsc machine by pushing the power button and it came back and now running fine.

 

(what I did)

The story is almost the same as the last two times (1st time, 2nd time).

I edited c1lsc.ini file using daqconfig and then shutdown daqd running fb.

Some indicators for c1lsc on the C1_FE_STATUS screen became red. So I hit the 'DAQ reload' button on the C1LSC_GDS_TP screen.

Then c1lsc died and didn't respond to ping.

  4705   Thu May 12 22:54:20 2011 ranaUpdateDAQInput Beam Naming change (no more IP)

 We decided to rename the Input Beam channels (while keeping temporary backwards compatible aliases) as:

C1:ASC-IB_POS_X, C1:ASC-IB_POS_Y, C1:ASC-IB_ANG_SUM, etc.

  4704   Thu May 12 18:21:19 2011 kiwamuUpdateIOOwelcome back IP_POS

IP_POS is back. 

I reconnected the cable to an interface card : D030238-A which has been labeled as "IP POS".

The card currently resides on the third crate rack from the top at the very right side in 1Y2. Also a rear side connection was modified a little bit.

I will clean up some cables because I tried some cables to see which was which.

Hopefully I will make a simple wiring diagram such that we will never forget the connections.

Quote from #4700

Quote from #4799

A cable for IP_POS has been disconnected at the LSC rack, 1Y2. Due to it currently no IP_POS signal shows up on the digital side.

I will check how it used to be and reconnect it.

 I found the disconnected cable, but I do not see the interface board at the LSC rack

  4703   Thu May 12 16:50:22 2011 SureshUpdateRF Systeminstallation of RF splitters in Demod boards

We have no plan to change the AS11 PD. I was referring to the AS11 Demod board which currently has the "Demodulator Preamp" circuit installed as a piggy back. In future I will append "_Demod" when I am referring to a demod board, to avoid confusion.

Quote:

Quote:

Rana also advised that we must use the boards which have the piggy-back amplifiers on those signals which are most useful.  We referred to Alberto's thesis and chose POY55 (MICH and SRCL), REFL11(PRCL)  and AS55 (DARM) as the most useful signals.  We currently have these amps on AS11, REFL11 and AS55.  We need to convert either AS11 or REFL11 into a POY55.  Since we need to troubleshoot REFL11, I thought we might as well modify that and in the process also fix its Q output.  So I renamed AS11 as REFL11 and will convert the old REFL11 into POY55 tomorrow. 

 I think we should leave them as is; the AS11 was made by taking into account the SB levels at the AS port and should not become REFL11. We should instead convert one of the old 25 or 33 MHz diodes into a POY55.

 

  4702   Thu May 12 10:27:02 2011 ranaUpdateRF Systeminstallation of RF splitters in Demod boards

Quote:

Rana also advised that we must use the boards which have the piggy-back amplifiers on those signals which are most useful.  We referred to Alberto's thesis and chose POY55 (MICH and SRCL), REFL11(PRCL)  and AS55 (DARM) as the most useful signals.  We currently have these amps on AS11, REFL11 and AS55.  We need to convert either AS11 or REFL11 into a POY55.  Since we need to troubleshoot REFL11, I thought we might as well modify that and in the process also fix its Q output.  So I renamed AS11 as REFL11 and will convert the old REFL11 into POY55 tomorrow. 

 I think we should leave them as is; the AS11 was made by taking into account the SB levels at the AS port and should not become REFL11. We should instead convert one of the old 25 or 33 MHz diodes into a POY55.

  4701   Thu May 12 09:00:23 2011 steveUpdateGeneraloptical table covers must be used

I found ITMX-ISCT and BS-ISCT ( IFO sensing-control table)  not covered this morning. Please do not leave tables open.

  4700   Thu May 12 08:54:11 2011 steveUpdateIOOIP_POS cable found

Quote:

A cable for IP_POS has been disconnected at the LSC rack, 1Y2. Due to it currently no IP_POS signal shows up on the digital side.

It looks like we disconnected the cable together with some unused cables when we were cleaning up the wiring of the LSC rack.

The cable, a shielded flat-cable, is supposed to send DC power to the QPD and send the signals from the QPD to an interface board on the LSC rack.

I will check how it used to be and reconnect it.

 I found the disconnected cable, but I do not see the interface board at the LSC rack

  4699   Thu May 12 05:25:26 2011 kiwamuUpdateIOOIP_POS disconnected

A cable for IP_POS has been disconnected at the LSC rack, 1Y2. Due to it currently no IP_POS signal shows up on the digital side.

It looks like we disconnected the cable together with some unused cables when we were cleaning up the wiring of the LSC rack.

The cable, a shielded flat-cable, is supposed to send DC power to the QPD and send the signals from the QPD to an interface board on the LSC rack.

I will check how it used to be and reconnect it.

  4698   Thu May 12 02:31:07 2011 SureshUpdateRF Systeminstallation of RF splitters in Demod boards

[Jamie, Koji, Suresh]

We replaced splitters in several demod boards as per the table given below:

 

Demod boards in which splitter has been replaced on 11May2011
Demod Board S. No. Power Splitter Frequency range of splitter (MHz) Phase unbalance from datasheet (deg) Amplitude Unbalance from datasheet (dB)
REFL55 118 PQW-2-90 30 to 90 90.14 0.92
AS11 021 PSCQ-2-51W 5 to 50 87.49 0.1
POY11 119 PSCQ-2-32 3.2 to 32 87.58 0.05
POY22 020 PSCQ-2-32 3.2 to 32 90.26 0.02
POY110 120 PSCQ--120 80 to 120 90.88 0.58

 

While doing a rough check of the boards I noticed that the REFL11 demod board had no signal on the Q output.

Rana also advised that we must use the boards which have the piggy-back amplifiers on those signals which are most useful.  We referred to Alberto's thesis and chose POY55 (MICH and SRCL), REFL11(PRCL)  and AS55 (DARM) as the most useful signals.  We currently have these amps on AS11, REFL11 and AS55.  We need to convert either AS11 or REFL11 into a POY55.  Since we need to troubleshoot REFL11, I thought we might as well modify that and in the process also fix its Q output.  So I renamed AS11 as REFL11 and will convert the old REFL11 into POY55 tomorrow.

Power splitter of different types have different pin-outs.  The way we mount a splitter depends on which type we are using.   I will detail the mounting scheme in a separate elog tomorrow.

 

 

 

  4697   Thu May 12 00:59:45 2011 ryan, ranaUpdatePSLReturn of the PSL temperature box

The PSL temperature box has returned to service, with some circuit modifications. The 1k resistors on all the temp. sensor inputs (R3, R4, R7, R8, R12, R12) were changed to 0 Ohm. Also, the 10k resistors R26, R28, R29, and R30 were changed to 10.2k metal film. The DCC document will be updated shortly. There is now an offset in the MINCOMEAS channel compared to the others, which will be corrected in the morning after looking at the overnight trend.

  4696   Wed May 11 23:02:52 2011 valeraUpdateASSDither angular stabilizitaion system update

This is what was done in past two days:

- The ETMY and ITMY pitch and yaw dofs are modulated at 40, 44, 42, 46 Hz respectively (oscillator A=30). The c1ass lockin numbers are 12, 14, 27, 29.

- The NAS55I signal is demodulated at the above frequencies. The demodulated I/Q signal phase is set to shift all signal into I-phase. The lockin inputs are bandpassed around respective frequency f with butter("Bandpass",2,f-0.5,f+0.5). The demod signals are then additionally low passed with butter ("Lowpass",4,0.5) so the servo ugf has to be below 0.5 Hz. The servo filter is p:z 0.0001:0.1.

- The ETMY demodulated signal is fed back to ITMY and visa versa.

- With the above 2x2 servo running we moved the input beam PZTs by hand to follow the cavity.

- At the end we offloaded the servo control signals to the SUS biases again by hand.

- The beam spot centering was estimated by unbalancing the ETMY/ITMY pitch/yaw coil combinations intentionally by 5%, which produces 1.3 mm shift of the node, and comparing the response to the residual signals.

- The dof set up currently is: ETMY pitch lockin 12 -> dof2, ITMY pitch lockin 14 -> dof4, ETMY yaw lockin 27 -> dof7, ITMY yaw lockin 29 -> dof9

- The next step is to demodulate the TRY(X) and servo the input beam PZTs

  4695   Wed May 11 22:54:53 2011 JenneUpdateTreasureETMY trans QPD installed

I put the ETMY trans QPD in. 

The ETMY trans beam was already going toward the TRY DC PD, and a CCD camera.  I put a beam splitter in that beam (reducing the power to TRY and the CCD by 50%), and sent my picked-off beam to the ETMY QPD.  Since there is a lens in this path to focus the beam onto TRY and the camera, I put the QPD ~the same distance from the lens as the camera.  Due to space requirements (because of all the green stuff on the table now), I had to put a Y1 turning mirror between the beam splitter and the QPD.  The beam is aligned onto the PD, although the signal isn't super strong.  When the PD is blocked, the sum is ~(-92 counts).  When the beam is on the PD, the sum is ~(-78 counts). 

  4694   Wed May 11 22:52:55 2011 kiwamuUpdateLSCstatus update and plan

Rana forced me to write this entry for summary because he didn't come to the 40m meeting today.

Status update :

    Interferometer Input Beam alignment with the PZTs.

     60 % done. The rest of the 40 % is to make the procedures automated.

     The beam spots on ITMY and ETMY are centered within ~1 mm accuracy.

     PRM, BS, ITMX, & ITMY actuator calibrations and PRC/MICH error calibrations

    Ongoing: First we will do it by hand, then some scripts will be made for the calibraion and resultant noise budget.

     F2A Suspension filter calculations.

      ETMY and ITMY are done. Need volunteer for ETMX, ITMX, BS, PRM, & SRM !!

      Bounce-Roll notch filters

      Leo is working on it. 25% complete...

     DC signal from RFPDs

      The RFPDs have a local SMA DC output as well as a DC output from their PD Interface cards in the LSC rack. We have hooked up some of the PD Interface DC outputs to the LSC ADCs but not tested??

Next steps:

  Installation of a temporary (Thorlabs) DCPD on either POY to see the intra-cavity power in the PRC. It would be ridiculous to put detectors on POX or POP since they're still clipped.

  D-phase and amplitude imbalance adjustment of the demod baords.

          make a script which uses pynds.

  Alignment of the full interferomter, starting from the X arm

  Loss measurements for the arms

  Schnupp asymmetry measurement

  

  4693   Wed May 11 21:58:30 2011 ranaUpdateSUSETMX_OPLEV : He-Ne replaced

Quote:

Quote:

I found that a He-Ne laser which has been used for ETMX_OPLEV was NOT giving the light.

Since I didn't find the switch key for it I have no idea if the laser is simply off or dead.

 The dead laser was replaced by new JDSU 1104P of  2.6mW. The return beam is big ~5 mm diameter of 0.3 mW,  1400 counts

 Whenever replacing any Oplev laser, please also put into the ELOG when it was installed so that we have an electronic record of the laser lifetime.

  4692   Wed May 11 17:20:24 2011 kiwamuConfigurationIOOloop diabled on PZT2

[Valera / Kiwamu]

The pointing of the incident beam to the interferometer has been jumping frequently.

Due to this jump the lock of the Y arm didn't stay for more than 2 min.

We turned off the strain gauge loop of PZT2-YAW and PZT2-PITCH, then the spot motion became solid and the Y arm locking became much more robust.

  4691   Wed May 11 17:10:04 2011 kiwamuConfigurationElectronicsfixed : MC3 LL PD has no signal

[Valera / Kiwamu]

It was because of a loose connection. Pushing the connector solved the issue.

looseconnectionMC3.png

We really have to think about making reliable connections and strain reliefs.

Quote from #4685

Yesterday we found that MC3 OSEM LL PD did not have a sensible signal - the readback was close to zero and it was making MC move around. I disabled the PD LL so that the damping is done with just three face plus side PDs. There still no signal from MC3 LL PD today. It needs debugging.

 

  4690   Wed May 11 16:04:36 2011 Larisa ThorneConfigurationElectronicsJenne-Seismometer LPF project

The schematic for the seismometer box from this last time  has been updated...

 

Koji was helpful for coming up with a general diagram for the voltage buffer amplifier, which has now been added to the configuration pictured below.

The only thing that remains now before I try to plot it with Eagle/LISO is to pick an op amp to use for the voltage buffer itself. Someone suggested THS4131 for that (upon Googling, it hit as a "high speed, low noise, fully-differential I/O amplifier"). It looks good, but is it the best option?

Attachment 1: STS2diagram2.png
STS2diagram2.png
  4689   Wed May 11 13:54:32 2011 steveUpdateLSCLSC power supplies

Aux- rack _1Y2 is just behind 1Y3  It contains Kepco Analoge DC power supplies for +- 5, 15 and 24V

Placing these power supplies away from the LSC rack was an effort to minimize pick up from them.

 

Attachment 1: P1070679.JPG
P1070679.JPG
  4688   Wed May 11 11:44:52 2011 kiwamuUpdateSUSf2p filters installed on ITMY

New f2p filters were installed on ITMY this morning. The statistical error of the coil gain setttigs are about 0.6 % at high frequency.

NEXT : PRM, SRM, BS, ITMX and ETMX.

 

f2p_itmy.png

 Pendulum mode = 0.988 Hz, Q = 1

C1:SUS-ITMY_ULCOIL_GAIN = 1.02482

C1:SUS-ITMY_URCOIL_GAIN = -1.06831

C1:SUS-ITMY_LLCOIL_GAIN = -0.996671

C1:SUS-ITMY_LRCOIL_GAIN = 0.91079

UL: fz = 1.014824 Hz, Q = 1.027150

UR: fz = 0.975038 Hz, Q = 0.98688

LL: fz = 1.000229 Hz, Q = 1.012378

LR: fz = 0.972688 Hz, Q = 0.946116

Quote from #4682

New f2p filters were installed on ETMY. 

The coil imbalances are now about 0.8% at high frequency (i.e. above the resonant freq of the pendulum mode)

  4687   Wed May 11 11:24:59 2011 steveUpdateSUSSRM-oplev qpd is back

Quote:

The SRM qpd cable was removed from the BS-table. It's path was changed from 1x4 to ITMY-table following the inner cable tray.

 Laser diode oplev SRM is working. Qpd matrix values were reset like others.

In 0.44mW, returning 0.1mW,  -500 counts.

Attachment 1: srm-prm_oplevs.jpg
srm-prm_oplevs.jpg
ELOG V3.1.3-